# EAGLE: Eigen Aggregation Learning for Object-Centric Unsupervised Semantic Segmentation

### v1

@@ Abstract
Semantic segmentation has innately relied on extensive pixel-level labeled annotated data, leading to the emergence of unsupervised methodologies. Among them, leveraging self-supervised Vision Transformers for unsupervised semantic segmentation (USS) has been making steady progress with expressive deep features. Yet, for semantically segmenting images with complex objects, a predominant challenge remains: the lack of explicit object-level semantic encoding in patch-level features. This technical limitation often leads to inadequate segmentation of complex objects with diverse structures. To address this gap, we present a novel approach, EAGLE, which emphasizes object-centric representation learning for unsupervised semantic segmentation. Specifically, we introduce EiCue, a spectral technique providing semantic and structural cues through an eigenbasis derived from the semantic similarity matrix of deep image features and color affinity from an image. Further, by incorporating our object-centric contrastive loss with EiCue, we guide our model to learn object-level representations with intra- and inter-image object-feature consistency, thereby enhancing semantic accuracy. Extensive experiments on COCO-Stuff, Cityscapes, and Potsdam-3 datasets demonstrate the state-of-the-art USS results of EAGLE with accurate and consistent semantic segmentation across complex scenes.
@@ 3.2.1 Adjacency Matrix Construction
(II) Semantic Similarity Matrix $\mathbf{A}_{\text{seg}}$: The semantic similarity matrix, denoted as $\mathbf{A}_{\text{seg}}\in\mathbb{R}^{N\times N}$, is formed by the product of tensor $\mathbf{S}$ and its transpose $\mathbf{S}^{\top}$. Tensor $\mathbf{S}$ is derived by hierarchically concatenating key attention features from the last three layers of a pretrained vision transformer, as processed through the segmentation head.
@@ 3.2.2 Eigendecomposition
To construct EiCue based on A, a Laplacian matrix is created. Formally, the Laplacian Matrix is expressed as $\mathbf{L}=\mathbf{D}-\mathbf{A}$, where $\mathbf{D}$ is the degree matrix of $\mathbf{A}$ defined as $\mathbf{D}(i,i)=\sum_{j=1}^{N}\mathbf{A}(i,j)$. In our approach, we utilize the normalized Laplacian matrix for its enhanced clustering capabilities. The symmetric normalized Laplacian matrix $\mathbf{L}_{\text{sym}}$ are defined as $\mathbf{L}_{\text{sym}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{L}\mathbf{D}^{-\frac{1}{2}}$. Then, via eigendecomposition on $\mathbf{L}_{\text{sym}}$, the eigenbasis $\mathbf{V}\in\mathbb{R}^{N\times N}$ is computed, where each column corresponds to a unique eigenvector. We then extract $k$ eigenvectors corresponding to the $k$ smallest eigenvalues and concatenate them into $\mathbf{\hat{V}}\in\mathbb{R}^{N\times k}$ where the $i^{th}$ row corresponds to the $k$ dimensional eigenfeature of the $i^{th}$ patch.
@@ 3.2.3 Differentiable Eigen Clustering
$\small\mathcal{L}_{\text{eig}}=-\frac{1}{N}\sum_{i=1}^{N}\bigg{(}\sum_{c=1}^{C}\Psi_{ic}\mathbf{P}_{ic}\bigg{)},$ (1)
[…]
@@ 3.3.1 Object-wise Prototypes
To extract the representative object level semantic features from projected feature $\mathbf{Z}$, we construct learnable prototypes $\Phi_{l}$ based on the object $l$ in aforementioned EiCue. As we describe next, semantically representative prototypes become the anchors for either pulling objects with similar semantics while pushing away the different ones.
@@ 3.3.2 Object-centric Contrastive Loss
$\small\mathcal{L}_{\text{nce}}^{x\leftrightarrow\tilde{x}}=\frac{1}{2}(\mathcal{L}_{\text{nce}}^{x\rightarrow\tilde{x}}+\mathcal{L}_{\text{nce}}^{\tilde{x}\rightarrow x}).$ (6)
@@ 4.1 Experimental Settings
Datasets. We evaluate on (1) COCO-Stuff [4], (2) Cityscapes [9], and (3) Potsdam-3 [22] datasets, in line with methodologies established in prior works [49, 16, 8, 22]. (1) The COCO-Stuff dataset composed of its detailed pixel-level annotations, facilitating comprehensive various object understanding, while (2) Cityscapes presents diverse urban street scenes. (3) The Potsdam-3 dataset is composed of satellite imagery. Following the class selection protocols from previous studies  [49, 16, 8, 22], we use 27 classes from both COCO-Stuff and Cityscapes. For Potsdam-3, we use all 3 classes (see supplement for result).
@@ B.3. Cityscapes
Additional qualitative results are available at Fig. 12.

### v3

@@ Abstract
Semantic segmentation has innately relied on extensive pixel-level annotated data, leading to the emergence of unsupervised methodologies. Among them, leveraging self-supervised Vision Transformers for unsupervised semantic segmentation (USS) has been making steady progress with expressive deep features. Yet, for semantically segmenting images with complex objects, a predominant challenge remains: the lack of explicit object-level semantic encoding in patch-level features. This technical limitation often leads to inadequate segmentation of complex objects with diverse structures. To address this gap, we present a novel approach, EAGLE, which emphasizes object-centric representation learning for unsupervised semantic segmentation. Specifically, we introduce EiCue, a spectral technique providing semantic and structural cues through an eigenbasis derived from the semantic similarity matrix of deep image features and color affinity from an image. Further, by incorporating our object-centric contrastive loss with EiCue, we guide our model to learn object-level representations with intra- and inter-image object-feature consistency, thereby enhancing semantic accuracy. Extensive experiments on COCO-Stuff, Cityscapes, and Potsdam-3 datasets demonstrate the state-of-the-art USS results of EAGLE with accurate and consistent semantic segmentation across complex scenes.
@@ 3.2.1 Adjacency Matrix Construction
(II) Semantic Similarity Matrix $\mathbf{A}_{\text{seg}}$: The semantic similarity matrix, denoted as $\mathbf{A}_{\text{seg}}\in\mathbb{R}^{N\times N}$, is formed by the product of tensor $\mathbf{S}$ and its transpose $\mathbf{S}^{\top}$. Tensor $\mathbf{S}$ is derived by hierarchically concatenating key attention features from the last three layers of a pretrained vision transformer, as processed through the segmentation head $\mathcal{S}_{\theta}$.
@@ 3.2.2 Eigendecomposition
To construct EiCue based on $\mathbf{A}$, a Laplacian matrix is created. Formally, the Laplacian Matrix is expressed as $\mathbf{L}=\mathbf{D}-\mathbf{A}$, where $\mathbf{D}$ is the degree matrix of $\mathbf{A}$ defined as $\mathbf{D}(i,i)=\sum_{j=1}^{N}\mathbf{A}(i,j)$. In our approach, we utilize the normalized Laplacian matrix for its enhanced clustering capabilities. The symmetric normalized Laplacian matrix $\mathbf{L}_{\text{sym}}$ are defined as $\mathbf{L}_{\text{sym}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{L}\mathbf{D}^{-\frac{1}{2}}$. Then, via eigendecomposition on $\mathbf{L}_{\text{sym}}$, the eigenbasis $\mathbf{V}\in\mathbb{R}^{N\times N}$ is computed, where each column corresponds to a unique eigenvector. We then extract $k$ eigenvectors corresponding to the $k$ smallest eigenvalues and concatenate them into $\mathbf{\hat{V}}\in\mathbb{R}^{N\times k}$ where the $i^{th}$ row corresponds to the $k$ dimensional eigenfeature of the $i^{th}$ patch.
@@ 3.2.3 Differentiable Eigen Clustering
$\small\mathcal{L}_{\text{eig}}^{\mathbf{x}}=-\frac{1}{N}\sum_{i=1}^{N}\bigg{(}\sum_{c=1}^{C}\Psi_{ic}\mathbf{P}_{ic}\bigg{)},$ (1)
where $C$ denotes pre-defined number of classes, $\Psi:=\textit{softmax}(\mathbf{P})$ and $\mathbf{P}_{ic}$ and $\Psi_{ic}$ represents the $i^{th}$ patch and the $c^{th}$ cluster number of $\mathbf{P}$ and $\Psi$. We apply same procedure to augmented image $\mathbf{\tilde{x}}$ to get $\mathcal{L}_{\text{eig}}^{\mathbf{\tilde{x}}}$. By minimizing $\mathcal{L}_{\text{eig}}=\frac{1}{2}(\mathcal{L}_{\text{eig}}^{\mathbf{x}}+\mathcal{L}_{\text{eig}}^{\mathbf{\tilde{x}}})$, we can obtain centers of clusters that enable more effective clustering. Then we obtain EiCue as
@@ 3.3.1 Object-wise Prototypes
To extract the representative object level semantic features from projected feature $\mathbf{Z}$, we construct adaptable prototypes $\Phi_{l}$ based on the object $l$ in aforementioned EiCue. As we describe next, semantically representative prototypes become the anchors for either pulling objects with similar semantics while pushing away the different ones.
@@
Thus, $\Phi_{l}$ acts as the semantic vector of object $l$, serving as an anchor for the following object-centric contrastive loss.
@@ 3.3.2 Object-centric Contrastive Loss
$\small\mathcal{L}_{\text{nce}}^{x\leftrightarrow\tilde{x}}=\mathcal{L}_{\text{nce}}^{x\rightarrow\tilde{x}}+\mathcal{L}_{\text{nce}}^{\tilde{x}\rightarrow x}.$ (6)
@@ 4.1 Experimental Settings
Datasets. We evaluate on (1) COCO-Stuff [4], (2) Cityscapes [9], and (3) Potsdam-3 [22] datasets, in line with methodologies established in prior works [49, 16, 8, 22]. (1) The COCO-Stuff dataset is composed of its detailed pixel-level annotations, facilitating comprehensive various object understanding, while (2) Cityscapes presents diverse urban street scenes. (3) The Potsdam-3 dataset is composed of satellite imagery. Following the class selection protocols from previous studies  [49, 16, 8, 22], we use 27 classes from both COCO-Stuff and Cityscapes. For Potsdam-3, we use all 3 classes (see supplement for result).
@@ B.3. Cityscapes
Additional qualitative results are available in Fig. 12.
@@ C.1.1 Ablation: Feature Type
C.1.2 Ablation: Prototype Selection Method
Table 6: The experimental results for the prototype selection method on the COCO-Stuff dataset.
[…]
Acc. mIoU Acc. mIoU
PCA 55.7 18.3 74.7 39.6
Centroid 59.0 24.9 75.8 42.1
Medoid 64.2 27.2 76.8 43.9
[…]
We further carry out ablation experiments comparing various methods for selecting prototypes. A prototype is a semantic vector that represents a single object, serving as an anchor that attracts semantic vectors within the object and repels the other ones. To select a semantic vector that represents an object, we can consider several options for choosing object-representative semantic vectors. (a) Principal Component Analysis (PCA): we can use PCA to find the direction of maximum variance in the data, and choose the vector that has the largest projection on the first principal component. (b) Centroid: calculates the mean vector of the set, where each component of the centroid is the average of that component across all vectors in the set. (c) Medoid: we can choose the vector that minimizes the sum of distances to all other vectors. This is less sensitive to outliers compared to the centroid. As reported through Table 6, we observed that leveraging the Medoid method demonstrated the best performance.
C.2. Additional Visualization of the Primary Elements of Eigen Aggregation Module
C.2.1 Eigenvectors
As illustrated in Fig. 14, we provide additional visualization of eigenvectors obtained from learnable feature $\mathbf{S}$ trained using COCO-Stuff dataset [4]. Our eigenvectors present remarkable capability in distinguishing objects while capturing within its object semantics.
C.2.2 EiCue
As shown in Fig. 15, we present additional visualizations that compare our EiCue model with the traditional K-means clustering approach. EiCue shows a significant improvement over clustering with K-means in identifying the semantic details of objects and discerning the comprehensive structure of images. This distinction is particularly evident in the way EiCue captures intricate object semantics and delineates the structural elements within the images. These observations substantiate our model’s claim that EiCue is proficient in recognizing object semantics and distinguishing structural components.
C.3. Application
Our EAGLE is designed for semantic segmentation, which predicts dense class prediction. Consequently, EAGLE is applicable to a variety of tasks that necessitate pixel-level semantic interpretation. A key technique in our approach is the eigendecomposition of the Laplacian matrix. Thus, we can use eigenvectors of images and these eigenvectors are able to capture the detailed semantic structures present in an image. As discussed in Section C.3.1, leveraging these eigenvectors allows us to precisely perform image matting and achieve localized image stylization.
C.3.1 Image Matting
Figure 9: We demonstrate our Laplacian matrix-based image matting. From left to right: the input image, eigenvector from our matrix, resultant image matte, CLIP-based segmentation [34], and edited image using our matte with existing text-driven image editing model [30]. For the editing in the last column, the text prompts used were plastic bag for the top image and white fur for the bottom image. Our matte offers clearer object boundaries than CLIPSeg, leading to superior editing quality.
The eigenvectors obtained through our proposed method are effective in distinguishing objects within images. To this end, our eigenvectors can be efficiently utilized for image matting and localized image stylization. The process of image matting involves the extraction of the foreground from an image, facilitating further manipulations like compositing onto a different background or selective editing. Historically, the matting task has been challenging due to intricate object edges and subtle transitions. The eigenvectors, adept at distinguishing distinct objects or features within images, provide a powerful solution to this challenge. Using our proposed method, we leverage the potential of these eigenvectors for refined image matting. As depicted in Fig. 9, the first column represents the input images. The subsequent column showcases the first eigenvector derived from our matrix, effectively highlighting the structure of the primary object. The third column portrays the resultant image matte, distinctly separating the object from its surroundings. In contrast, the fourth column, representing the CLIP-based segmentation, while reasonable, but fails to provide as delicate boundaries as our eigenvector-based technique. A notable difference can be observed in the image of the boat. Our method adeptly separates the pillar of the boat, whereas the CLIP-based approach fails to isolate it, erroneously including the trees in the background as part of the foreground. The final column presents the edited images, emphasizing the utility of our matte for selective edits, ensuring that the distinguished object can be stylized without affecting the background.
D. Implementation Details
In this section, we discuss details of correspondence distillation loss (Section D.1.), model architecture (Section D.2.), and hyperparameters (Section D.3).
D.1. Correspondence Distillation Loss
By employing a correspondence distillation loss [16], we enhanced the stability of the training process by ensuring reliable graph Laplacian initialization. The original correspondence distillation loss is defined as
$\mathcal{L}_{\text{cd}}(\check{\mathbf{F}},\check{\mathbf{S}},b)=-\sum\left(\check{\mathbf{F}}-b\right)\check{\mathbf{S}},$ (8)
where $\check{\mathbf{F}}$ and $\check{\mathbf{S}}$ is computed as a cosine distance using $\mathbf{F}$ and $\mathbf{S}$. Here, $\mathbf{F}$ is a feature obtained from the activation map at the final layer of the vision transformer with a given image and $\mathbf{S}$ is the projection of $\mathbf{F}$ using segmentation head $\mathcal{S}_{\theta}$. Since we leverage attention keys in place of $\mathbf{F}$, we substitute $\mathbf{F}$ with $\mathbf{K}$ and revise $\mathbf{S}$ to be $\mathbf{S}=\mathcal{S}_{\theta}(\mathbf{K})$. Within the framework of existing correspondence distillation loss [16], which involves three distinct loss functions, our method modifies and utilizes two of these components: (a) the augmented image correspondence distillation loss and the (b) random image correspondence distillation loss.
Although Eq. (8) is applied to both types of loss, the difference lies in what each correspondence tensor represents. (a) In the augmented image correspondence distillation loss, $\mathbf{\check{K}}_{\text{aug}}$ and $\mathbf{\check{S}}_{\text{aug}}$ is computed as a cosine distance between $\mathbf{K},\mathbf{\tilde{K}}$ and $\mathbf{S},\mathbf{\tilde{S}}$, respectively. $\mathbf{\tilde{K}}$ and $\mathbf{\tilde{S}}$ are the results for the $\mathbf{\tilde{x}}$, which is the augmented images of $\mathbf{x}$, created through the same aforementioned process. While (b) in the random image correspondence distillation loss, $\mathbf{\check{K}}_{\text{rand}}$ and $\mathbf{\check{S}}_{\text{rand}}$ is computed as a cosine distance between $\mathbf{K},\mathbf{\acute{K}}$ and $\mathbf{S},\mathbf{\acute{S}}$, respectively. $\mathbf{\acute{K}}$ and $\mathbf{\acute{S}}$ are the results for the random images from the entire dataset, created through the same aforementioned process. In the Eq. (8), $b$ is defined as the shift of the feature value and remained fixed throughout the training process. In contrast, we modified $b_{\text{aug}}$ and $b_{\text{rand}}$ to dynamically adapt based on the $\check{\mathbf{K}}$ and $\check{\mathbf{S}}$ in both losses, where $b_{\text{aug}}$ represents the $b$ in augmented image correspondence distillation loss, and $b_{\text{rand}}$ represents the $b$ in random image correspondence distillation loss. Here is the formula we used:
$\small b_{\text{aug}}=\left\lvert{\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}\check{\mathbf{K}}_{ij}}-{\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}\check{\mathbf{S}}_{ij}-k_{\text{shift}}}\right\rvert,$ (9)
$b_{\text{rand}}=\left({\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}\check{\mathbf{K}}_{ij}}+{\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}\check{\mathbf{S}}_{ij}}-k_{\text{shift}}\right)\times v_{\text{shift}},$ (10)
where $H$ and $W$ refer to the height and width of the feature tensor, respectively. Here, $k_{\text{shift}}$ and $v_{\text{shift}}$ are determined as hyperparameters (see Section D.3.) Thus, the final correspondence distillation loss that we use is defined as
$\small\mathcal{L}_{\text{corr}}=\mathcal{L}_{\text{cd}}(\check{\mathbf{K}}_{\text{aug}},\check{\mathbf{S}}_{\text{aug}},b_{\text{aug}})+\mathcal{L}_{\text{cd}}(\check{\mathbf{K}}_{\text{rand}},\check{\mathbf{S}}_{\text{rand}},b_{\text{rand}}),$ (11)
which is the summation of augmented image correspondence distillation loss and random image correspondence distillation loss.
D.2. Detailed Architecture
Figure 10: Detailed architecture of segmentation head and projection head used in our method.
Image Encoder. For all experiments, we basically leverage DINO [6] pretrained ViT-S/8, ViT-S/16 and ViT-B/8 [14] as an image encoder. Specifically, we initialize ViT with a teacher weight of DINO. As mentioned before, we extract attention keys hierarchically from ViT and then concatenate them into a single feature tensor $\mathbf{K}$ (for details, see Section 3.1 in the main manuscript). Then, we apply channel-wise dropout ($p=0.1$) to feature tensor $\mathbf{K}$ before feeding to segmentation head $\mathcal{S}_{\theta}$.
Segmentation Head. We illustrate the detailed architecture of the segmentation head and projection head in Fig. 10. For a fair performance comparison with existing models, we employ the same approach for the segmentation head $\mathcal{S}_{\theta}$ as used in the previous models [16, 49]. This non-linear segmentation head $\mathcal{S}_{\theta}$ consists of simple linear layers. The input is a tensor $\mathbf{K}$ with a dimension of $D_{K}$. This tensor first passes through a linear layer that transforms its dimension from $D_{K}$ to $D_{S}$, where $D_{S}$ represents the desired dimension for the output of $\mathcal{S}_{\theta}$. Following the initial linear transformation, there is a ReLU (Rectified Linear Unit) activation function, which introduces non-linearity to the process. The output of the activation layer is then fed into another linear layer, which once again maps the dimension from $D_{K}$ to $D_{S}$. The outputs of the two pathways are then combined via a summation operation. The summation consolidates the linearly transformed input and the non-linearly transformed input. The result is the tensor $\mathbf{S}$ with the dimension $D_{S}$, which is the output of the segmentation head.
Projection Head. As shown in Fig. 10, we project semantic tensor $\mathbf{S}$ to $\mathbf{Z}$ to facilitate object-centric contrastive learning. The basic concept of projection head $\mathcal{Z}_{\xi}$ is to project the tensor without transforming its input dimension. Following this concept, we form a projection head with a single linear layer that maps its dimension from $D_{S}$ to $D_{Z}$. Here, we note that while we use different notations $D_{S}$ and $D_{Z}$ for clarity and ease of explanation, the actual dimensions represented by these notations are the same as $D_{S}=D_{Z}$.
D.3. Hyperparameters
Table 7: Hyperparameters used in EAGLE. LR refers learning rate.
[…]
$\lambda_{\text{obj}}$ 0.3 0.3
$\lambda_{\text{sc}}$ 0.7 0.7
$\lambda_{\text{nce}}$ 0.9 0.7
$k_{\text{shift}}$ 0 0.11
$v_{\text{shift}}$ 3.5 3.5
step 200 380
LR $\mathcal{S}_{\theta},\mathcal{Z}_{\xi},\Phi$ 0.0005 0.0005
$\mathbf{C}$ 0.00005 0.0004
[…]
In this section, we carefully describe hyperparameters in Table D.3. Hyperparameters that are used throughout our series of experiments. In the table above, “step” refers to the number of training iterations required for the $\lambda_{\text{nce}}$ to increase from 0 to the number indicated in $\lambda_{\text{nce}}$ row. “LR” indicates learning rate and $\mathcal{S}_{\theta},\mathcal{Z}_{\xi},\Phi$ shares same learning rate.
E. Discussion
E.1. Failure Cases
Figure 11: Failure cases of EAGLE.
Unsupervised semantic segmentation (USS), unlike the fully-supervised approach, is quite challenging as it predicts classes for each pixel without labeled data. While USS is likely to show much better performance in the future, our model represents a step in its evolution and thus comes with certain limitations. Fig. 11 illustrates the failure cases of EAGLE, which has been trained to capture object-level semantics. The first to the third column is from the COCO-Stuff dataset and the remaining columns are from the Cityscapes dataset. In the first column, our model fails to segment objects properly. This is due to the narrow color distribution of the input image and the limited variety of object semantics present in the image, leading to a failure in creating a high-quality adjacency matrix for EiCue. In the second and third columns, we observed that our results successfully implemented object-level semantics but made errors in matching the object class. Within column four, we see our results that accurately segment and correctly classify car, tree, building, and person, but incorrectly categorize gravel paths as a different class instead of road. Similarly, in the last column, there were no critical errors for objects other than sidewalk. However, even when viewed with the human eye, the input image presents a challenging scenario in distinguishing between road and sidewalk.
E.2. Future Works
Throughout our manuscript, we demonstrated that leveraging EiCue through graph Laplacian effectively captures the semantic structure of an image. However, constructing an adjacency matrix and forming a Laplacian matrix entails a relatively high computational cost. This approach does not affect the inference time in our framework, but it does require more training time compared to using solely deep-based methods. In our research, we compute the adjacency matrix for every feature vector of an image. While EAGLE shows state-of-the-art results, regarding every single feature vector is not the most computationally efficient, suggesting that improvements to EiCue could be made by sampling only vital features based on other knowledge of the image and constructing the Laplacian matrix accordingly. Additionally, as our primary focus is on object-level semantics, this approach may not be directly applicable to domains like medical imaging. Therefore, it is crucial to engage in research that uncovers knowledge about object-level semantics, which is applicable across multiple domains and holds significant potential for widespread use in the field of computer vision.
Figure 12: Additional qualitative results of Cityscapes dataset [9] trained with ViT-B/8 backbone.
Figure 13: Additional qualitative results of COCO-Stuff dataset [4] trained with ViT-S/8 backbone.
Figure 14: Additional visualization of the eigenvector obtained by training the COCO-Stuff dataset [4] with ViT-S/8 as the backbone.
Figure 15: Additional comparison between K-means and our EiCue, with EiCue demonstrating enhanced performance in discerning object semantics and structures relative to K-means.